OpenStudy (anonymous):

A store had 235 MP3 players in the month of January. Every month, 30% of the MP3 players were sold and 50 new MP3 players were stocked in the store. Which recursive function best represents the number of MP3 players in the store f(n) after n months? f(n) = 0.7 × f(n - 1) + 50, f(0) = 235, n > 0 f(n) = 237 - 0.7 × f(n - 1) + 50, f(0) = 235, n > 0 f(n) = 0.3 × f(n - 1) + 50, f(0) = 235, n > 0 f(n) = 237 + 0.7 × f(n - 1) + 50, f(0) = 235, n > 0

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

@ math essentials

OpenStudy (anonymous):

f(n) = 0.7(n-1) + 50, f(0) = 235, n>0, showing 30% sold and 50 added

OpenStudy (anonymous):

(05.06) Nick works two jobs to pay for college. He tutors for $15 per hour and also works as a bag boy for $8 per hour. Due to his class and study schedule, Nick is only able to work up to 20 hours per week, but must earn at least $150 per week. If t represents the number of hours Nick tutors and b represents the number of hours he works as a bag boy, which system of inequalities represents this scenario?

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

15 t + 8 b >= 150 t + b <= 20 if you wanted to solve this, you would turn these into equalities

OpenStudy (anonymous):

Can you look at the pic please?

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

a. would require him to work too many hours, b and c. make no sense, "less than or greater than" no wonder you were puzzled. Only d {none of these] applies.

OpenStudy (anonymous):

Yes lol thank you

OpenStudy (anonymous):

Glad to help.

OpenStudy (anonymous):

@douglaswinslowcooper